# Formalizing zeta and L-functions in Lean

### v1

@@ Abstract.
The Riemann zeta function, and more generally the $L$-functions of Dirichlet characters, are among the central objects of study in analytic number theory. We report on a project to formalize the theory of these objects in Lean’s mathlib library.
@@ 1.1. Mathematical background
Dirichlet introduced these functions in order to study how the primes are distributed among the residue classes modulo $N$; they played a central role in his proof that for any $a$ with $\gcd(a,N)=1$, there exist infinitely many primes $p$ with $p=a\bmod N$.
@@ 1.2. Results formalized in this project
There are analogous results for Dirichlet characters, including the non-vanishing on $\{s:\Re(s)\geqslant 1\}$. From this we deduce Dirichlet’s theorem on primes in arithmetic progressions as Nat.setOf_prime_and_eq_mod_infinite.
We hope that a proof of the prime number theorem – and, more generally, of a “quantitative” version of Dirichlet’s theorem, giving an asymptotic for the number of primes $\leqslant X$ in a given congruence class – will be formalized in Lean in the near future; this is the goal of the “PrimeNumberTheorem+” project led by Alex Kontorovich and Terry Tao [PNT+].
@@ 1.3. Related work
  • Manuel Eberl and his collaborators have formalized a substantial part of the theory of Riemann zeta and $L$-functions in the “Isabelle” theorem prover [Eb19]. Building on this, Song and Yao [SY24] recently announced an Isabelle formalisation of the prime number theorem with the “classical” error term $O(x\exp(-C\sqrt{\log x}))$).
    While the main results of this Isabelle-based project are comparable to (and indeed go considerably further than) the results of the Lean formalisation described here, there are major differences in methodology, particularly in the proof of the functional equation. For this result, Riemann gave two proofs in [Ri59]. The Isabelle proof is based on contour integrals, while we follow the theta-function proof. As we shall see below, this involved the formalisation of a range of interesting results along the way, and has many natural generalisations, such as in the theory of modular forms.
@@ 2. Implementing $L$-series in mathlib
The number-theoretic significance of the Riemann zeta function, Dirichlet $L$-functions and numerous other $L$-functions is related to the sequence of coefficients in their expansion as Dirichlet series and its properties (e.g., via an expression as an Euler product). So an implementation of Dirichlet series (which we will call $L$-series below) and their relevant properties is a prerequisite for applications like the Prime Number Theorem or Dirichlet’s Theorem. At the beginning of this project, a rudimentary implementation of $L$-series in mathlib existed, which associated a function $\mathbb{C}\to\mathbb{C}$ to an “arithmetic function” $f$; arithmetic functions are implemented as functions $f\colon\mathbb{N}\to R$ with $f(0)=0$ (here $R$ is some semiring, which needs to have a coercion to $\mathbb{C}$ in our context). This made it possible to write LSeries $\zeta$ or LSeries $\mu$ for the Dirichlet series associated to the arithmetic functions $\zeta$ (taking the value $1$ everywhere outside zero) or the Möbius function $\mu$; but, for example, LSeries $\chi$ with a Dirichlet character $\chi$ or LSeries $\Lambda$ with the von Mangoldt function $\Lambda$ did not type-check because there are no coercions from DirichletCharacter $\mathbb{C}$ $n$ or from ArithmeticFunction $\mathbb{R}$ to ArithmeticFunction $\mathbb{C}$. It turned out that it was not possible to set up a general mechanism of coercion from ArithmeticFunction $R$ to ArithmeticFunction $R^{\prime}$ when there is a coercion from $R$ to $R^{\prime}$, and hard-coding the relevant cases $\mathbb{N}\hookrightarrow\mathbb{Z}\hookrightarrow\mathbb{R}\hookrightarrow\mathbb{C}$ led to a considerable amount of code duplication.
To avoid that, we changed the implementation of $L$-series to not use the type of arithmetic functions. A first attempt using the argument type $\mathbb{N}+\to\mathbb{C}$ (where $\mathbb{N}+$ is the subtype of $\mathbb{N}$ consisting of positive natural numbers) was abandoned in favour of the approach described below because it turned out that the necessity of going back and forth between $\mathbb{N}$ and $\mathbb{N}+$ was causing some friction; also, some API for $\mathbb{N}+$ was missing.
@@
It turns out that with this set-up, the formalisation of operations with and properties of $L$-series proceeds quite smoothly. This includes, for example, the Euler product representation of Dirichlet $L$-series.
@@ 3.1. Fourier analysis
Note that in this formulation we have decay assumptions on both $f$ and $\hat{f}$. This is easy to verify in our application, where both $f$ and $\hat{f}$ are Gaussian functions. However, in other applications it is advantageous to formulate all the hypotheses in terms of $f$ alone. Building on the contributions described above, S. Gouëzel has recently added to mathlib a proof that the Fourier transform of a Schwartz function is also a Schwartz function, from which it follows that the Poisson summation formula holds for any Schwartz function
(SchwartzMap.tsum_eq_tsum_fourierIntegral).
@@ 3.2. Jacobi’s theta function
In order to extend this to Dirichlet character $L$-series, it is necessary to consider the more general two-variable theta function
@@
This satisfies a functional equation generalizing that of the one-variable function (and proved by the same method); and its Mellin transform gives the analytic continuation and functional equation of the even Hurwitz zeta function (hurwitzZetaEven)
$\zeta_{\alpha}^{\text{ev}}(s)=\sum_{\begin{subarray}{c}n\in\mathbb{Z}\\ n+\alpha\neq 0\end{subarray}}\frac{1}{|n+\alpha|^{s}},\qquad\alpha\in\mathbb{R}.$
@@
To include the case of odd Dirichlet characters, we need to study yet another variant of the theta-series, defined by
@@
This is related to the odd Hurwitz zeta function hurwitzZetaOdd, defined by
$\zeta_{\alpha}^{\text{odd}}(s)=\sum_{\begin{subarray}{c}n\in\mathbb{Z}\\ n+\alpha\neq 0\end{subarray}}\frac{\operatorname{sign}(n+\alpha)}{|n+\alpha|^{s}},$
@@
In the Mathlib implementation of the above results, the definition and properties of the Jacobi theta function is approximately 900 code lines, and the proof of the analytic continuation and functional equation of Dirichlet $L$-functions another 3300 lines. (This surprisingly large total is in part because several intermediate steps, e.g. justifying the interchange of integration and summation required to relate $\int_{0}^{\infty}t^{s-1}\frac{\theta(it)-1}{2}\,\mathrm{d}t$ to $\sum_{n}n^{-s}$ when the latter converges, were done in considerably greater generality than absolutely required for this project, in order to enable later generalisations; see below.)
@@
Since the above results on Poisson summation and Jacobi theta functions were added to Mathlib, they have found a surprising and unexpected real-world application: the transformation formula for theta functions was used in implementing verified algorithms for privacy-preserving computation for Amazon Web Services [dM+04].
3.3. Generalisations
@@
Hence we have set up these arguments in an axiomatic framework which is intended to apply to any modular form $L$-function, following the arguments in [DS05, Chapter 5]. We introduce a concept we call an “FE-pair”: this is a pair of locally integrable functions $f,g$ on $\mathbb{R}_{\geqslant 0}$ such that
@@
We show that for any FE-pair, the Mellin transforms of $f$ and $g$ have meromorphic continuation (with poles determined by the asymptotics of $f$ and $g$) and satisfy a functional equation relating values at $s$ and $k-s$.
The basic example of an FE-pair is $f(x)=g(x)=\theta(ix)$ (with $k=\tfrac{1}{2}$); and the two-variable theta functions also give rise to FE-pairs, again with $k=\tfrac{1}{2}$. The formalisation of modular form theory in mathlib is ongoing; and we hope to use the FE-pair framework to formalize properties of $L$-functions of higher weight modular forms in future projects.
@@ 4. Special values of $L$-functions
Lastly, we describe the formalization of Euler’s “Basel problem” formula $\zeta(2)=\tfrac{\pi^{2}}{6}$, and its generalisations. The proof we formalised for these also uses Fourier analysis, although in a rather different fashion from above. It arises by studying the “periodized Bernoulli function” periodizedBernoulli, i.e. the unique function on $\mathbb{R}/\mathbb{Z}$ whose restriction to $[0,1)$ is the $k$-th Bernoulli polynomial $B_{k}$. A standard computation shows that the $n$-th Fourier coefficient of this function is $\tfrac{-k!}{(2\pi in)^{k}}$. For even $k\geqslant 2$, considering the pointwise sum at 00, and applying the results on Fourier-series convergence described above, gives the standard formula for $\zeta(2k)$ in terms of the $2k$-th Bernoulli number $B_{2k}(0)$ (riemannZeta_two_mul_nat). This includes the Basel problem as a special case.
This approach generalises immediately to computing values of Hurwitz zeta functions. The result is most tidily stated in terms of values at negative integers,
$\zeta_{\alpha}(1-k)=\frac{-B_{k}(\alpha)}{k}\quad\text{for $\alpha\in[0,1]$ and $k\geqslant 2$}$
(HurwitzZeta.hurwitzZeta_neg_nat).
A disadvantage of our approach is that we must exclude the case $k=1$. In this case the above formula remains valid; but our proof does not, since the Fourier series concerned has $c_{n}(f)=-1/(2\pi in)$ and hence does not satisfy the condition $\sum_{n\in\mathbb{Z}}|c_{n}(f)|<\infty$. For instance, the Gregory–Leibniz formula $1-\tfrac{1}{3}+\tfrac{1}{5}-\tfrac{1}{7}+\dots=\tfrac{\pi}{4}$, giving the value at $s=1$ of the $L$-series of the unique nontrivial Dirichlet character modulo 4, is not a special case of our results [^3] This particular formula is already in Mathlib as Real.tendsto_sum_pi_div_four, but the method used does not seem to generalise to other Dirichlet characters.. Including this case would require extending Mathlib’s Fourier theory library to treat conditionally-convergent Fourier series, which would be an interesting project for the future.
@@ 5. Pitfalls: totalizing functions
As pointed out by K. Buzzard (personal communication), there is an issue if $s=1$. Our construction assigns some value to the expression riemannZeta 1; but it is far from obvious what this value is. If this junk value happened to be zero, then the conjecture we had formulated would not be the Riemann Hypothesis; it would be trivially false!
@@
Remark: Such oversights might have unfortunate consequences in the context of efforts to automate discovery of proofs via artificial intelligence. In this situation, mistakes in formalizing the statements of conjectures might lead to mathematicians believing that a conjecture had been proved (or disproved), when in fact the system had proved or disproved a superficially similar but much easier statement, “exploiting” an oversight in formulating the conjecture in some mathematically uninteresting corner case.
@@ 6. Applications to primes in arithmetic progressions
One key ingredient that is needed here is a proof that the Dirichlet $L$-functions donot vanish on the closed right half-plane $\Re(s)\geqslant 1$. When $s\neq 1$ or the Dirichlet character is not quadratic, this follows fairly easily from the nonnegativity of a certain trigonometric polynomial. The remaining case, $L(\chi,1)\neq 0$ when $\chi$ is a quadratic Dirichlet character, requires a more subtle argument. We establish that an $L$-series with nonnegative coefficients (and positive coefficient at $1$) that extends to an entire function takes positive values on all of $\mathbb{R}$ (LSeries.positive_of_differentiable_of_eqOn). Assuming that $L(\chi,1)=0$, we obtain that $L(\chi,s)\zeta(s)$ extends to an entire function (the zero of $L(\chi,{\cdot})$ cancels the pole of $\zeta$) and is an $L$-series with nonnegative coefficients, so it must be positive in particular at $s=-2$. But $\zeta(-2)=0$, leading to a contradiction. This establishes the general non-vanishing result, DirichletCharacter.LFunction_ne_zero_of_one_le_re.
To prove Dirichlet’s Theorem, we consider the $L$-series associated to the function that sends $n$ to $\Lambda(n)$ (where $\Lambda$ is the von Mangoldt function) when $n\equiv a\bmod q$ and to zero otherwise. This $L$-series is a linear combination of negative logarithmic derivatives of $L$-series associated to the Dirichlet characters mod $q$: LSeries_residueClass_eq. This step in the proof uses the orthogonality relations for (Dirichlet) characters, which we also contributed to mathlib. The non-vanishing result above then implies that these logarithmic derivatives extend continuously to $\Re(s)\geqslant 1$, except for a simple pole at $s=1$ when $a$ and $q$ are coprime. The existence of this simple pole then fairly quickly leads to a proof that $\sum_{n\equiv a\bmod q}\Lambda(n)/n$ diverges, which in turn easily implies the existence of infinitely many prime numbers $p\equiv a\bmod q$. We state this final result in two versions, Nat.setOf_prime_and_eq_mod_infinite (the set of these primes is infinite) and Nat.forall_exists_prime_gt_and_eq_mod (for every natural number $n$, there exists a larger prime $p\equiv a\bmod q$).

### v3

@@ Abstract.
The Riemann zeta function, and more generally the $L$-functions of Dirichlet characters, are among the central objects of study in analytic number theory. We report on a project to formalize the theory of these objects in Lean’s ‘Mathlib’ library, including a proof of Dirichlet’s theorem on primes in arithmetic progressions and a formal statement of the Riemann hypothesis.
@@ 1.1. Mathematical background
Dirichlet introduced these functions in order to study how the primes are distributed among the residue classes modulo $N$; they played a central role in his proof that for any $a$ with $\gcd(a,N)=1$, there exist infinitely many primes $p$ with $p\equiv a\bmod N$.
@@ 1.2. Results formalized in this project
There are analogous results for Dirichlet characters, including the non-vanishing on $\{s:\Re(s)\geqslant 1\}$. From this we deduce Nat.setOf_prime_and_eq_mod_infinite, Dirichlet’s theorem on primes in arithmetic progressions.
@@ 1.3. Related work
  • Manuel Eberl and his collaborators have formalized a substantial part of the theory of Riemann zeta and $L$-functions in the “Isabelle” theorem prover [Eb19]. Building on this, Song and Yao [SY24] recently announced an Isabelle formalization of the prime number theorem with the “classical” error term $O(x\exp(-C\sqrt{\log x}))$).
    While the main results of this Isabelle-based project are comparable to (and indeed go considerably further than) the results of the Lean formalization described here, there are major differences in methodology, particularly in the proof of the functional equation. For this result, Riemann gave two proofs in [Ri59]. The Isabelle proof is based on contour integrals, while we follow the theta-function proof. As we shall see below, this involved the formalization of a range of interesting results along the way, and has many natural generalizations, such as in the theory of modular forms.
  • Building on the formalizations described in this paper, the “PrimeNumberTheorem+” project led by Alex Kontorovich and Terry Tao [PNT+] has recently formalized in Lean a proof of the Prime Number Theorem (via the Wiener–Ikehara theorem), which will hopefully be merged into Mathlib in the near future. Future goals of the PNT+ project include formalizing an explicit bound on the error term, and extending the results to counting primes in arithmetic progressions, strengthening Dirichlet’s theorem by giving an asymptotic for the number of primes $\leqslant X$ in a given congruence class.
@@ 2. Implementing $L$-series in Mathlib
The number-theoretic significance of the Riemann zeta function, Dirichlet $L$-functions and numerous other $L$-functions is related to the sequence of coefficients in their expansion as Dirichlet series and its properties (e.g., via an expression as an Euler product). So an implementation of Dirichlet series (which we will call $L$-series below) and their relevant properties is a prerequisite for applications like the Prime Number Theorem or Dirichlet’s Theorem.
At the beginning of this project, a rudimentary implementation of $L$-series in Mathlib existed, which associated a function $\mathbb{C}\to\mathbb{C}$ to an “arithmetic function” $f$; arithmetic functions are implemented as functions $f\colon\mathbb{N}\to R$ with $f(0)=0$ (here $R$ is some semiring, which needs to have a coercion to $\mathbb{C}$ in our context). This made it possible to write LSeries $\zeta$ or LSeries $\mu$ for the Dirichlet series associated to the arithmetic functions $\zeta$ (taking the value $1$ everywhere outside zero) or the Möbius function $\mu$; but, for example, LSeries $\chi$ with a Dirichlet character $\chi$, or LSeries $\Lambda$ with the von Mangoldt function $\Lambda$, did not type-check because there are no coercions from DirichletCharacter $\mathbb{C}$ $n$ or from ArithmeticFunction $\mathbb{R}$ to ArithmeticFunction $\mathbb{C}$.
It turned out that it was not easily possible to set up a general coercion from ArithmeticFunction $R$ to ArithmeticFunction $R^{\prime}$ when there is a coercion from $R$ to $R^{\prime}$, because the necessary coercion instance would have too many free parameters; and hard-coding the relevant cases $\mathbb{N}\hookrightarrow\mathbb{Z}\hookrightarrow\mathbb{R}\hookrightarrow\mathbb{C}$ led to a considerable amount of code duplication. See the discussion on the Lean Zulip chat [^2] https://leanprover.zulipchat.com/#narrow/channel/144837-PR-reviews/topic/.2310725.20and.20.2310728.3A.20L-series/near/422590170 for details on this and also on how we arrived at the current design, which is described below.
To avoid these problems, we changed the implementation of $L$-series to not use the type of arithmetic functions. A first attempt using the argument type $\mathbb{N}+\to\mathbb{C}$ (where $\mathbb{N}+$ is the subtype of $\mathbb{N}$ consisting of positive natural numbers) was abandoned in favour of the approach described below because it turned out that the necessity of going back and forth between $\mathbb{N}$ and $\mathbb{N}+$ was causing some friction; also, some API for $\mathbb{N}+$ was missing.
@@
It turns out that with this set-up, the formalization of operations with and properties of $L$-series proceeds quite smoothly. This includes, for example, the Euler product representation of Dirichlet $L$-series.
@@ 3.1. Fourier analysis
Note that in this formulation we have decay assumptions on both $f$ and $\hat{f}$. This is easy to verify in our application, where both $f$ and $\hat{f}$ are Gaussian functions. However, in other applications it is advantageous to formulate all the hypotheses in terms of $f$ alone. Building on the contributions described above, S. Gouëzel has recently added to Mathlib a proof that the Fourier transform of a Schwartz function is also a Schwartz function, from which it follows that the Poisson summation formula holds for any Schwartz function (SchwartzMap.tsum_eq_tsum_fourierIntegral).
@@ 3.2. Jacobi’s theta function
In order to extend this to Dirichlet character $L$-series, it is necessary to consider the more general two-variable theta function (formalized as jacobiTheta2)
@@
This satisfies a functional equation generalizing that of the one-variable function (and proved by the same method); and its Mellin transform gives the analytic continuation and functional equation of the even Hurwitz zeta function (HurwitzZeta.hurwitzZetaEven)
$\zeta_{\alpha}^{\text{ev}}(s)=\tfrac{1}{2}\sum_{\begin{subarray}{c}n\in\mathbb{Z}\\ n+\alpha\neq 0\end{subarray}}\frac{1}{|n+\alpha|^{s}},\qquad\alpha\in\mathbb{R}.$
@@
[…]
@@
This is related to the odd Hurwitz zeta function (HurwitzZeta.hurwitzZetaOdd), defined by
$\zeta_{\alpha}^{\text{odd}}(s)=\tfrac{1}{2}\sum_{\begin{subarray}{c}n\in\mathbb{Z}\\ n+\alpha\neq 0\end{subarray}}\frac{\operatorname{sign}(n+\alpha)}{|n+\alpha|^{s}},$
@@
In the Mathlib implementation of the above results, the definition and properties of the Jacobi theta function is approximately 900 code lines, and the proof of the analytic continuation and functional equation of Dirichlet $L$-functions another 3300 lines. (This surprisingly large total is in part because several intermediate steps, e.g. justifying the interchange of integration and summation required to relate $\int_{0}^{\infty}t^{s-1}\frac{\theta(it)-1}{2}\,\mathrm{d}t$ to $\sum_{n}n^{-s}$ when the latter converges, were done in considerably greater generality than absolutely required for this project, in order to enable later generalizations; see below.)
@@
Since the above results on Poisson summation and Jacobi theta functions were added to Mathlib, they have found a surprising and unexpected real-world application: the transformation formula for theta functions was used in implementing verified algorithms for privacy-preserving computation for Amazon Web Services [dM+24].
3.3. Generalizations
@@
Hence we have set up these arguments in an axiomatic framework which is intended to apply to any modular form $L$-function, following the arguments in [DS05, Chapter 5]. We introduce a concept we call a “functional equation pair”, or “FE-pair” for short: this is a pair of locally integrable functions $f,g$ on $\mathbb{R}_{\geqslant 0}$ such that
@@
The above definition is formalized as WeakFEPair (with StrongFEPair being the special case where the constant terms of $f,g$ at $\infty$ are both zero). We show that for any FE-pair, the Mellin transforms of $f$ and $g$ have meromorphic continuation (with poles and residues determined by the asymptotics of $f$ and $g$) and satisfy a functional equation relating values at $s$ and $k-s$ (WeakFEPair.functional_equation).
The basic example of an FE-pair is $f(x)=g(x)=\theta(ix)$ (with $k=\tfrac{1}{2}$); and the two-variable theta functions also give rise to FE-pairs, again with $k=\tfrac{1}{2}$. The formalization of modular form theory in Mathlib is ongoing; and we hope to use the FE-pair framework to formalize properties of $L$-functions of higher weight modular forms in future projects.
@@ 4. Special values of $L$-functions
Lastly, we describe the formalization of Euler’s “Basel problem” formula $\zeta(2)=\tfrac{\pi^{2}}{6}$, and its generalizations. The proof we formalized for these also uses Fourier analysis, although in a rather different fashion from above. It arises by studying the “periodized Bernoulli function” periodizedBernoulli, i.e. the unique function on $\mathbb{R}/\mathbb{Z}$ whose restriction to $[0,1)$ is the $k$-th Bernoulli polynomial $B_{k}(x)$. A standard computation shows that the $n$-th Fourier coefficient of this function is $\tfrac{-k!}{(2\pi in)^{k}}$. For even $k\geqslant 2$, considering the pointwise sum at 00, and applying the results on Fourier-series convergence described above, gives the standard formula for $\zeta(2k)$ in terms of the $2k$-th Bernoulli number $B_{2k}(0)$ (riemannZeta_two_mul_nat). This includes the Basel problem as a special case.
This approach generalizes immediately to computing values of Hurwitz zeta functions. The result is most tidily stated in terms of values at negative integers (HurwitzZeta.hurwitzZeta_neg_nat):
$\zeta_{\alpha}(-k)=\frac{-B_{k+1}(\alpha)}{k+1}\quad\text{for $\alpha\in[0,1]$and $k\geqslant 1$}.$
A disadvantage of our approach is that we must exclude the case $k=0$. In this case the above formula remains valid for $\alpha\in(0,1)$; but our proof no longer applies, since the Fourier series concerned has $c_{n}(f)=-1/(2\pi in)$ and hence does not satisfy the condition $\sum_{n\in\mathbb{Z}}|c_{n}(f)|<\infty$. For instance, the Gregory–Leibniz formula $1-\tfrac{1}{3}+\tfrac{1}{5}-\tfrac{1}{7}+\dots=\tfrac{\pi}{4}$, giving the value at $s=1$ of the $L$-series of the unique nontrivial Dirichlet character modulo 4, is not a special case of our results [^4] This particular formula is already in Mathlib as Real.tendsto_sum_pi_div_four, but the method used does not seem to generalize to other Dirichlet characters. Since the original draft of this paper was written, X. Roblot has contributed to Mathlib a proof of Dedekind’s class number formula for general number fields, NumberField.tendsto_sub_one_mul_dedekindZeta_nhdsGT; applying this to quadratic fields, it should be relatively straightforward to deduce formulae for $L(\chi,1)$ for any quadratic Dirichlet character.. Including this case would require extending Mathlib’s Fourier theory library by adding more general criteria for the Fourier series of a periodic function to converge to the function (such as the Dirichlet–Jordan criterion for functions of bounded variation); this would be an interesting project for the future.
@@ 5. Pitfalls: totalizing functions
As pointed out by K. Buzzard and J. Ellenberg (in discussions following a Twitter post by Buzzard about this work), there is an issue if $s=1$. Our construction assigns some value to the expression riemannZeta 1; but it is far from obvious what this value is. If this junk value happened to be zero, then the conjecture we had formulated would not be the Riemann Hypothesis; it would be trivially false!
@@
Remark
Such oversights might have unfortunate consequences in the context of efforts to automate discovery of proofs via artificial intelligence. In this situation, mistakes in formalizing the statements of conjectures might lead to mathematicians believing that a conjecture had been proved (or disproved), when in fact the system had proved or disproved a superficially similar but much easier statement, “exploiting” an oversight in formulating the conjecture in some mathematically uninteresting corner case.
@@ 6. Applications to primes in arithmetic progressions
One key ingredient that is needed here is a proof that the Dirichlet $L$-functions do not vanish on the closed right half-plane $\Re(s)\geqslant 1$. When $s\neq 1$ or the Dirichlet character is not quadratic, this follows fairly easily from the Euler product representation and the non-negativity of a certain trigonometric polynomial. The remaining case, $L(\chi,1)\neq 0$ when $\chi$ is a quadratic Dirichlet character, requires a more subtle argument. We establish that if an $L$-series with non-negative real coefficients (and positive coefficient at $1$) converges in some half-plane, and has analytic continuation as an entire function, then its analytic continuation must take positive real values on all of $\mathbb{R}$ (LSeries.positive_of_differentiable_of_eqOn). Assuming that $L(\chi,1)=0$, we obtain that $L(\chi,s)\zeta(s)$ extends to an entire function (the zero of $L(\chi,{\cdot})$ cancels the pole of $\zeta$) and is an $L$-series with nonnegative coefficients, so it must be positive in particular at $s=-2$. But $\zeta(-2)=0$, leading to a contradiction. This establishes the general non-vanishing result, DirichletCharacter.LFunction_ne_zero_of_one_le_re.
To prove Dirichlet’s Theorem, we consider the $L$-series associated to the function that sends $n$ to $\Lambda(n)$ (where $\Lambda$ is the von Mangoldt function) when $n\equiv a\bmod q$ and to zero otherwise. This $L$-series is a linear combination of logarithmic derivatives of $L$-series associated to the Dirichlet characters mod $q$: ArithmeticFunction.vonMangoldt.LSeries_residueClass_eq. This step in the proof uses the orthogonality relations for (Dirichlet) characters, which we also contributed to Mathlib. The non-vanishing result above then implies that these logarithmic derivatives extend continuously to $\Re(s)\geqslant 1$, except for a simple pole at $s=1$ when $a$ and $q$ are coprime. The existence of this simple pole then fairly quickly leads to a proof that $\sum_{n\equiv a\bmod q}\Lambda(n)/n$ diverges, which in turn easily implies the existence of infinitely many prime numbers $p\equiv a\bmod q$. We state this final result in two versions, Nat.setOf_prime_and_eq_mod_infinite (the set of these primes is infinite) and Nat.forall_exists_prime_gt_and_eq_mod (for every natural number $n$, there exists a larger prime $p\equiv a\bmod q$).